# Demystifying Group Relative Policy Optimization: Its Policy Gradient is a U-Statistic

### 03 Mar 2026 v2

@@ 4 Main results
  • Policy optimization: Section 4.2 shifts the focus to the sub-optimality gap of the learned policy. In particular, Lemma 6 presents a finite-sample upper bound on this gap, building upon which Theorem 7 establishes a scaling law that offers insights into the optimal choice of the group size. Theorem 8 further establishes the consistency of the parameter estimates and the asymptotic distribution of the sub-optimality gap without assuming parameter identifiability. These results verify the oracle property of the GRPO policy (Corollary 9) as well as its optimality (Corollary 10).
  • Practical considerations: Section LABEL:sec:ext of the Supplementary Material extends the aforementioned analyses to accommodate (i) reward standardization; (ii) importance sampling and (iii) the KL divergence penalty. This section establishes the connection between the gradient estimator and U-statistics (Lemma LABEL:lem:pgobjective), characterizes the gradient estimator’s MSE (Theorem LABEL:thm:practicalgradMSE), and derives the policy learning target (Proposition LABEL:prop:target).
@@ 4.1 Group relative gradient evaluation
$g(x;\theta)=\mathbb{E}^{\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(Y\mid x)(Z|X=x)];$
@@
where $\text{trace}[\bullet]$ denote the trace of a matrix, $\Sigma_{\text{oracle}}(x;\theta)$ denotes the asymptotic covariance matrix of the oracle gradient estimator $G\text{Cov}(\widehat{g}_{\text{oracle}}(x;\theta))$ and $\|\bullet\|$ denotes the $\ell_{2}$-norm of a vector.
@@
In the minibatch setting, the MSE is further reduced by a factor of $B^{-1}$. The first term in the second line of (9) represents the inherent variance arising from prompt sampling, which is independent of $G$ and shared by all gradient estimators (e.g., vanilla, oracle). Equation (9) is instrumental in deriving our scaling law in the next section. To elaborate, suppose we fix the sampling budget per prompt as $N=B\times G$. Then $B=N/G$, so the first term scales linearly with $G$, the second term is constant when $N$ is fixed, and the third term scales inversely with $G$. Consequently, there exists an optimal choice of $G$ that is neither too small nor too large, balancing the first and third terms to minimize the MSE. As we will show later, the MSE is closely related to the suboptimality gap of the learned policy, and thus there exists a corresponding $G$ that minimizes this suboptimality gap.
@@ 4.2 Group relative policy optimization
By definition, a smaller gap indicates a policy closer to the optimal one. Additionally, the suboptimality gap measures only the quality of the final policy obtained at the last iteration. In the RL literature, an alternative metric is regret (e.g., Jaksch et al., 2010), which quantifies the cumulative difference between the optimal policy and the sequence of intermediate policies across iterations. However, we adopt the suboptimality gap because it better reflects practical LLM applications: once training is complete, only the final policy is deployed in practice. This criterion also aligns with the evaluation metric in deployment-efficient RL (e.g., Huang et al., 2022).
@@
The sequence of learning rates $\{\eta_{i}\}_{i\geq 1}$ in Algorithm 1 satisfies either (a) a constant schedule where $\eta_{i}=\beta$ for some $\beta<(2L)^{-1}$, or (b) an $1/i$ schedule where $\eta_{i}=i^{-1}\beta$ for some constant $\beta>(2\mu)^{-1}$.
@@
Finally, Assumption 5 considers two standard learning rate schedules. The constant schedule (a) is common in practical applications (Sheng et al., 2025). To the contrary, the $1/i$ schedule (b) is motivated by stochastic approximation theory: it ensures that the suboptimality gap theoretically achieves the optimal convergence rate (Kushner and Yin, 2003).
@@
Suppose Assumptions 1 and 35 hold. Suppose each $C_{i}^{(b,g)}$ in Algorithm 1 is either a function of $X^{(b)}$ or a GRPO-type baseline in (4). Let $M=\sup_{i\geq 0}\textrm{MSE}(\widehat{g}(\theta_{i}))$ denote the uniform upper bound on the MSE of this algorithm’s gradient estimator. Then for any $n\geq 1$, under the constant schedule (a), the output policy of $\pi_{\theta_{n}}$ satisfies
$\Delta(\pi_{\theta_{n}})\leq(1-2\mu\beta+L\mu\beta^{2})^{n}\Delta(\pi_{\theta_{0}})+\frac{L\beta^{2}M}{4\mu\beta-2L\mu\beta^{2}}.$
Under the $1/i$ schedule (b), we have
$\Delta(\pi_{\theta_{n}})\leq\max\Big(\frac{(1+\epsilon)L\beta^{2}M}{(4\mu\beta-2)n},\frac{c}{n}\Big),$
@@
In GRPO, sampling reasoning traces from the learning policy (Line 4 of Algorithm 1) is computationally expensive. With a fixed sampling budget, Equation (11) reveals a trade-off in the allocation of computational resources. A small $G$ allows for a larger batch size $B$ or more iterations $n$, which reduces the first variance term in (11) or accelerates the convergence rates established in Lemma 6. Conversely, a large $G$ reduces the higher-order residual term (the third term in (11)).
Our scaling law proposed in (12) explicitly balances this trade-off. Crucially, the optimal group size $G^{*}$ is universal: it is independent of the budget $N$, the number of iterations $n$ and the learning rate schedule. Instead, $G^{*}$ depends solely on the underlying data generating process and the geometry of the policy space. This makes our result highly practical for implementation, as the optimal group size remains constant for a given task regardless of the total compute available. We will verify this observation empirically in Section 5. While the constants $c_{i}$ are formally defined as suprema over the parameter space, they may yield conservative theoretical bounds. In practice, these suprema can be relaxed by estimating the constants at a few representative parameter values (e.g., those of the reference model).
@@
We say the estimator $\theta_{n}$ is consistent if $d(\theta_{n},\Theta^{*})\stackrel{{\scriptstyle P}}{{\to}}0$ as $n\to\infty$. To address the second challenge, we recognize that while the parameter estimates are non-identifiable, the suboptimality gap remains identifiable. We thus shift our analysis to the asymptotic distribution of the suboptimality gap rather than the parameters themselves.
We consider the asymptotic regime where the number of iterations $n\to\infty$, while holding other parameters such as the batch size $B$ and group size $G$ constant. For this analysis, we adopt the $1/i$ schedule for the learning rate. This is motivated by Lemma 6: under the constant learning rate schedule, the sub-optimality gap might not converge to zero when $B$ and $G$ are fixed. While one could theoretically allow $B$ and $G$ to grow with $n$ or explore alternative decay schedules, these configurations would introduce substantial technical complexity into the proofs without altering our major findings. We focus on this specific regime to more clearly elucidate the performance of GRPO.
@@
$\Theta^{*}$ is a connected set. The Hessian matrices at $\Theta^{*}$ all share the same rank $r$. Moreover, there exists some orthogonal projection matrix $Q$ and a strictly negative definite matrix $H$ such that $Q^{\top}Q=I_{r}$ and $Q^{\top}H(\theta^{*})Q=H$ for all $\theta^{*}\in\Theta^{*}$.
@@
There exists some matrix $\Gamma$ such that
$\mathbb{E}\left\{(\widehat{g}(\theta_{i})-g(\theta_{i}))(\widehat{g}(\theta_{i})-g(\theta_{i}))^{\top}\Big|\theta_{i}\right\}\to\Gamma,$
almost surely as $i\to\infty$.
@@
Suppose Assumptions 7 and 8 hold additionally. Then the output policy satisfies
@@
where $r$ is the rank of the Hessian matrix, $\{\chi_{1,k}^{2}\}_{k=1}^{r}$ are i.i.d. $\chi_{1}^{2}$ random variables, and $\{w_{k}\}_{k=1}^{r}$ are positive weights determined by the eigenvalues of the asymptotic covariance matrix of the gradient estimators (see the Supplementary Material).
Assumption 6 is mild and commonly imposed in the literature (e.g., Schmidt-Hieber, 2020). The first part of Assumption 7 relaxes the standard identifiability condition by allowing the population parameter to form a connected manifold rather than a unique point. The second part assumes that, after projection onto a suitable $r$-dimensional subspace, the parameter becomes identifiable with a strictly negative definite Hessian. This effectively decomposes the parameter space into an identifiable component captured by $Q$ and a non-identifiable component along its orthogonal complement. Assumption 8 requires the conditional covariance of the gradient estimator to converge almost surely. While potentially strong, this assumption is necessary to establish the asymptotic normality. Without it, the limiting distribution may not exist.
@@
Suppose the assumptions in Theorem 8 hold. Then GRPO’s weights satisfy $BG(w_{k,\rm GRPO}-w_{k,\rm oracle})\to 0$ as $G\to\infty$. Consequently, the suboptimality gap of GRPO is asymptotically equivalent to that of the oracle algorithm.
@@
Suppose the assumptions in Theorem 8 hold. Suppose Assumption 2 further holds. Then for any meta-algorithm whose baseline term is a function of the prompt only, its weights satisfy $w_{k}\geq w_{k,\rm oracle}+O(B^{-1}G^{-2})$. Consequently, as $G\to\infty$, its suboptimality gap is asymptotically no smaller than that of GRPO.
@@ 5.1 Oracle property in gradient evaluation
We first evaluate the MSE of the three gradient estimators introduced in Section 4.1: vanilla, GRPO-type, and oracle. To conduct this comparison, we construct a synthetic arithmetic dataset consisting of 500 questions. Each question is a medium-difficulty integer arithmetic problem, sampled uniformly from five categories: (i) two-step addition or subtraction, (ii) three-step addition or subtraction, (iii) single-step multiplication, (iv) integer division, and (v) addition or subtraction with parentheses. For each generated problem, we record the ground-truth solution. We then query an LLM (the prompt is provided in Section LABEL:app:prompt of the Supplementary Material), extract the solution from its output, and compute a binary reward indicating whether the solution equals the ground truth.
We evaluate gradients aggregated over the 500 questions (see Equation (8)) with respect to three target policies, derived from three models with progressively stronger reasoning capabilities: (i) Qwen/Qwen2.5-0.5B Base model, (ii) Qwen/Qwen2.5-0.5B Instruct model, and (iii) Qwen/Qwen2.5-0.5B Instruct model with in-context learning (ICL, Brown et al., 2020). For the third model, in addition to the prompt, we provide a few-shot in-context demonstration containing several arithmetic questions with solutions (see Section LABEL:app:prompt of the Supplementary Material). These models produce target policies of different quality: the base model is expected to achieve the lowest accuracy, while the ICL model attains the highest. We also consider multiple groups sizes $G\in\{4,8,16,32,64\}$ to estimate the gradient and use Monte Carlo simulations to evaluate each gradient estimator’s MSE and report their associated $95\%$ confidence interval in Figure 4.
@@
    The oracle estimator (green line) achieves the smallest MSE, as expected. Nevertheless, with a moderately large group size ($G=8$), the MSE of the GRPO-type estimator is already close to that of the oracle estimator. As $G$ increases further (e.g., $G=32$ or $64$), the two estimators become nearly indistinguishable, confirming the oracle property in Corollary 4.
@@ 5.2 Optimal group size for policy optimization
As shown in our theoretical analysis and empirical results (Section 5.2), increasing the group size $G$ reduces the MSE of the GRPO-type gradient estimator. However, a larger $G$ also increases computational cost, as more outputs must be sampled per prompt. Under a fixed sampling budget (i.e., a fixed total number of sampled outputs), our scaling law in Theorem 7 shows that the optimal group size $G^{*}$ depends only on the data and the policy model, and is universal with respect to other parameters such as the number of iterations $n$ and the total budget per prompt $N$. In this section, we verify this universality using two widely adopted math reasoning benchmark datasets, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
@@
We next use the MATH dataset to assess the universality of the optimal group size $G^{*}$ across different sampling budgets $N$. The procedure closely follows that used for GSM8K: we consider the same candidate group sizes $G\in\{4,8,16,32,64,128\}$ and train the model separately for each choice of $G$. The differences are that, for MATH, we fine-tune a larger and more powerful Qwen2.5-Math-7B model, and we evaluate three different sampling budgets, $N\in\{1024,2048,4096\}$. Test accuracies of the resulting trained models are reported in Table 2.

### 22 Mar 2026 v3

@@ 4 Main results
  • Policy optimization: Section 4.2 shifts the focus to the sub-optimality gap of the learned policy. In particular, Lemma 6 presents a finite-sample upper bound on this gap, upon which Theorem 7 establishes a scaling law that offers insights into the optimal choice of the group size. Theorem 8 further establishes the consistency of the parameter estimates and derives the asymptotic distribution of the sub-optimality gap as a mixture of $\chi^{2}$ random variables without assuming parameter identifiability. These results verify the oracle property of the GRPO policy (Corollary 9) as well as its optimality (Corollary 10).
  • Practical considerations: Section A of the Supplementary Material extends the aforementioned analyses to accommodate (i) reward standardization; (ii) importance sampling and (iii) the KL divergence penalty. This section connects the gradient estimator with U-statistics (Lemma 11), characterizes the gradient estimator’s MSE (Theorem 12), and derives the consistency of the parameter estimates (Proposition 13).
@@ 4.1 Group relative gradient evaluation
$g(x;\theta)=\mathbb{E}^{\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(Y\mid x)\mathbb{E}(Z|X=x)];$
@@
where $\text{trace}[\bullet]$ denotes the trace of a matrix, $\Sigma_{\text{oracle}}(x;\theta)$ denotes the asymptotic covariance matrix of the oracle gradient estimator $G\text{Cov}(\widehat{g}_{\text{oracle}}(x;\theta))$ and $\|\bullet\|$ denotes the $\ell_{2}$-norm of a vector.
@@
In the minibatch setting, the MSE is further reduced by a factor of $B^{-1}$. The first term in the second line of (9) represents the inherent variance arising from prompt sampling, which is independent of $G$ and shared by all gradient estimators (e.g., vanilla, oracle). Equation (9) is instrumental in deriving our scaling law detailed in the next section. To elaborate, suppose we fix the sampling budget per prompt as $N=B\times G$. Then $B=N/G$, so the first term scales linearly with $G$, the second term is constant when NN is fixed, and the third term scales inversely with $G$. Consequently, there exists an optimal choice of $G$ that is neither too small nor too large, balancing the first and third terms to minimize the MSE. As we will show later, the MSE is closely related to the suboptimality gap of the learned policy, and thus there exists a corresponding $G$ that minimizes this suboptimality gap.
@@ 4.2 Group relative policy optimization
By definition, a smaller gap indicates a policy closer to the optimal one. Additionally, the suboptimality gap measures only the quality of the final policy obtained at the last iteration. In the RL literature, an alternative metric is regret, which quantifies the cumulative difference between the optimal policy and the sequence of intermediate policies across iterations. However, we adopt the suboptimality gap because it better reflects practical LLM applications: once training is complete, only the final policy is deployed in practice. This criterion also aligns with the evaluation metric in deployment-efficient RL (e.g., Huang et al., 2022).
@@
Additionally, the score function $\nabla_{\theta}\log\pi_{\theta}$ is uniformly bounded and continuous in $\theta$.
@@
The sequence of learning rates $\{\eta_{i}\}_{i\geq 1}$ in Algorithm 1 satisfies either (a) a constant schedule where $\eta_{i}=\beta$ for some $0<\beta<(2L)^{-1}$, or (b) an $1/i$ schedule where $\eta_{i}=i^{-1}\beta$ for some constant $\beta>(2\mu)^{-1}$.
@@
Finally, Assumption 5 considers two standard learning rate schedules. The constant schedule (a) is common in practical applications (Sheng et al., 2025). To the contrary, the $1/i$ schedule (b) is motivated by stochastic approximation theory: it ensures that the suboptimality gap theoretically achieves the optimal convergence rate (Zhang, 2016).
@@
Suppose Assumptions 1 and 35 hold. Suppose each $C_{i}^{(b,g)}$ in Algorithm 1 is either a function of $X^{(b)}$ or a GRPO-type baseline in (4). Let $M=\sup_{\theta}\textrm{MSE}(\widehat{g}(\theta))$ denote the uniform upper bound on the MSE of this algorithm’s gradient estimator. Then for any $n\geq 1$, under the constant schedule (a), the output policy of $\pi_{\theta_{n}}$ satisfies
$\mathbb{E}[\Delta(\pi_{\theta_{n}})]\leq(1-2\mu\beta+L\mu\beta^{2})^{n}\mathbb{E}[\Delta(\pi_{\theta_{0}})]+\frac{L\beta^{2}M}{4\mu\beta-2L\mu\beta^{2}},$
where the expectation is taken with respect to the randomness in the estimator $\theta_{n}$ and any randomness in the initial parameter $\theta_{0}$. Under the $1/i$ schedule (b), we have
$\mathbb{E}[\Delta(\pi_{\theta_{n}})]\leq\max\Big(\frac{(1+\epsilon)L\beta^{2}M}{(4\mu\beta-2)n},\frac{c}{n}\Big),$
@@
In GRPO, sampling reasoning traces from the learning policy (Line 4 of Algorithm 1) is computationally expensive. With a fixed sampling budget, Equation (11) reveals a trade-off in the allocation of computational resources. A small $G$ allows for a larger batch size $B$ and more iterations $n$, which reduces the first variance term in (11) and accelerates the convergence rates established in Lemma 6. Conversely, a large $G$ reduces the higher-order residual term (the third term in (11)).
Our scaling law proposed in (12) explicitly balances this trade-off. Crucially, the optimal group size $G^{*}$ is universal: it is independent of the budget NN, the number of iterations $n$ and the learning rate schedule. Instead, $G^{*}$ depends solely on the underlying data generating process and the geometry of the policy space. This makes our result highly practical for implementation, as the optimal group size remains constant for a given task regardless of the total compute available. We will verify this observation empirically in Section 5. While the constants $c_{i}$ are formally defined as suprema over the parameter space, they may yield conservative theoretical bounds. In practice, these suprema can be relaxed by estimating the constants at a few representative parameter values (e.g., those of the reference model).
@@
We say the estimator $\theta_{n}$ is consistent if $d(\theta_{n},\Theta^{*})\stackrel{{\scriptstyle P}}{{\to}}0$ as $n\to\infty$. To address the second challenge, we recognize that while the parameter estimates are non-identifiable, the suboptimality gap remains identifiable. We thus analyze the asymptotic distribution of the suboptimality gap rather than the parameters themselves.
Consider the asymptotic regime where the number of iterations $n\to\infty$ and other parameters such as the batch size $B$ and group size $G$ are constant. For this analysis, we adopt the $1/i$ schedule for the learning rate. This is motivated by Lemma 6: under the constant learning rate schedule, the sub-optimality gap might not converge to zero when $B$ and $G$ are fixed. While one could theoretically allow $B$ and $G$ to grow with $n$ or explore alternative decay schedules, these configurations introduce substantial technical complexity into the proofs without altering our major findings. We focus on this specific regime to more clearly elucidate the performance of GRPO, and impose the following conditions.
@@
$\Theta^{*}$ is closed and convex, and all Hessian matrices at $\Theta^{*}$ share the same rank $r$. Moreover, there exists an orthogonal projection matrix $Q$ and a strictly negative definite matrix $H^{*}$ such that $Q^{\top}Q=I_{r}$ and $Q^{\top}H(\theta^{*})Q=H^{*},\forall\theta^{*}\in\Theta^{*}$.
@@
There exists a (possibly random) matrix $\Gamma$ such that the following holds almost surely as $n\to\infty$,
$\mathbb{E}\left\{(\widehat{g}(\theta_{n})-g(\theta_{n}))(\widehat{g}(\theta_{n})-g(\theta_{n}))^{\top}\Big|\theta_{n}\right\}\to\Gamma.$
Assumption 9 (Weak strong concavity).
Let $\Pi_{\Theta^{*}}$ denote the projection operator such that $\Pi_{\Theta^{*}}(\theta):=\arg\min_{x\in\Theta^{*}}d(\theta,x)$, we have
$\langle g(\theta),\Pi_{\Theta^{*}}(\theta)-\theta\rangle\geq\mu d^{2}(\theta,\Theta^{*}),\quad\forall\theta\in\Theta.$
Assumption 6 is mild and commonly imposed in the literature (e.g., Schmidt-Hieber, 2020). Assumption 7 relaxes the standard identifiability condition by allowing the population parameter to form a connected manifold rather than a unique point. It assumes that, after projection onto a suitable $r$-dimensional subspace, the parameter becomes identifiable with a strictly negative definite Hessian $H^{*}$. This effectively decomposes the parameter space into an identifiable component defined through the transformation matrix $Q$ and a non-identifiable component along its orthogonal complement. Assumption 8 is not overly restrictive, as the limiting matrix $\Gamma$ is allowed to be random. The resulting asymptotic distribution can be interpreted as a mixture distribution: first, $\Gamma$ is realized, and then the limiting distribution is obtained conditionally on $\Gamma$. Finally, Assumption 9 imposes a weak strong concavity (WSC) condition. It is weaker than strong concavity (which entails a unique optimizer) but stronger than PL (Assumption 4). In contrast to strong concavity, WSC does not require the objective function to be globally concave in all directions, but only along directions that point toward the optimal set $\Theta^{*}$. Thus, the objective function $\mathbb{E}^{\pi_{\theta}}(Z)$ may remain flat along other directions, allowing multiple optimizers to exist. Under the $L$-smoothness condition (Assumption 3), it can be shown that WSC implies the PL condition.
@@
Suppose Assumptions 79 hold additionally. Then the output policy satisfies
@@
where $r$ is the rank of the Hessian matrix, $\{\chi_{1,k}^{2}\}_{k=1}^{r}$ are i.i.d. $\chi_{1}^{2}$ random variables, and $\{w_{k}\}_{k=1}^{r}$ are positive weights arranged in non-increasing order, determined by the asymptotic covariance matrix of the gradient estimators (see step (iv) of the proof of Theorem 8 in the Supplementary Material).
@@
Suppose the assumptions in Theorem 8 hold. Suppose the projected covariance matrix $\Omega(\theta^{*})=\textrm{Cov}(Q^{\top}\widehat{g}(\theta^{*}))$ depends on $\theta^{*}\in\Theta^{*}$ only through $Q^{\top}\theta^{*}$ and is non-singular. Then GRPO’s weights and oracle weights satisfy $w_{k,\rm GRPO}-w_{k,\rm oracle}=O(G^{-2})$. Consequently, the suboptimality gap of GRPO is asymptotically equivalent to that of the oracle algorithm as $G\to\infty$.
@@
Suppose the assumptions in Corollary 9 hold. Suppose $Q^{\top}\nabla_{\theta}\log\pi_{\theta^{*}}(Y|X)[\nabla_{\theta}\log\pi_{\theta^{*}}(Y|X)]^{\top}Q$ is conditionally uncorrelated with $Z$ given $X$, for some $\theta^{*}\in\Theta^{*}$. Then for any meta-algorithm whose baseline term is a function of the prompt only, its weights satisfy $w_{k}\geq w_{k,\rm oracle}+O(G^{-2})$. Consequently, as $G\to\infty$, its suboptimality gap is asymptotically no smaller than that of GRPO.
@@ 5.1 Oracle property in gradient evaluation
We first evaluate the MSE of the three gradient estimators introduced in Section 4.1: vanilla, GRPO-type, and oracle. To conduct this comparison, we construct a synthetic arithmetic dataset consisting of 500 questions. Each question is a medium-difficulty integer arithmetic problem, sampled uniformly from five categories: (i) two-step addition or subtraction, (ii) three-step addition or subtraction, (iii) single-step multiplication, (iv) integer division, and (v) addition or subtraction with parentheses. For each generated problem, we record the ground-truth solution. We then query an LLM (the prompt is provided in Section C of the Supplementary Material), extract the solution from its output, and compute a binary reward indicating whether the solution equals the ground truth.
We evaluate gradients aggregated over the 500 questions (see Equation (8)) with respect to three target policies, derived from three models with progressively stronger reasoning capabilities: (i) Qwen/Qwen2.5-0.5B Base model, (ii) Qwen/Qwen2.5-0.5B Instruct model, and (iii) Qwen/Qwen2.5-0.5B Instruct model with in-context learning (ICL, Brown et al., 2020). For the third model, in addition to the prompt, we provide a few-shot in-context demonstration containing several arithmetic questions with solutions (see Section C of the Supplementary Material). These models produce target policies of different quality: the base model is expected to achieve the lowest accuracy, while the ICL model attains the highest. We also consider multiple groups sizes $G\in\{4,8,16,32,64\}$ to estimate the gradient and use Monte Carlo simulations to evaluate each gradient estimator’s MSE and report their associated $95\%$ confidence interval in Figure 4.
@@
    The oracle estimator (green line) achieves the smallest MSE, as expected. Nevertheless, with a moderately large group size ($G=8$), the MSE of the GRPO-type estimator is already close to that of the oracle estimator. As $G$ increases further (e.g., $G=32$ or 6464), the two estimators become nearly indistinguishable, confirming the oracle property in Corollary 4.
@@ 5.2 Optimal group size for policy optimization
As shown in our theoretical analysis and empirical results (Section 5.2), increasing the group size $G$ reduces the MSE of the GRPO-type gradient estimator. However, a larger $G$ also increases computational cost, as more outputs must be sampled per prompt. Under a fixed sampling budget (i.e., a fixed total number of sampled outputs), our scaling law in Theorem 7 shows that the optimal group size $G^{*}$ depends only on the data and the policy model, and is universal with respect to other parameters such as the number of iterations $n$ and the total budget per prompt NN. In this section, we verify this universality using two widely adopted math reasoning benchmark datasets, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
@@
We next use the MATH dataset to assess the universality of the optimal group size $G^{*}$ across different sampling budgets NN. The procedure closely follows that used for GSM8K: we consider the same candidate group sizes $G\in\{4,8,16,32,64,128\}$ and train the model separately for each choice of $G$. The differences are that, for MATH, we fine-tune a larger and more powerful Qwen2.5-Math-7B model, and we evaluate three different sampling budgets, $N\in\{1024,2048,4096\}$. Test accuracies of the resulting trained models are reported in Table 2.